OpenStudy (anonymous):

A net force of 25N is applied for 5.7s to a 12kg box initially at rest. What is the speed of the box at the end of the 5.7s interval a)12m/s b)1.8m/s c)30m/s d)3.0m/s e)7.5m/s Please explain as I need help on learning how to do this

OpenStudy (snowsurf):

You will need Newton Second law \[F =ma\]And you will need a kinematic equation to get final velocity after acceleration has been applied. We are going to assume acceleration does not vary so acceleration is constant. So we can use this equation\[Vf = Vi + at\] We know F/m = a and t is time in seconds. Vf = Vi + at. \[V_{f} = V _{i} + at\] But since a =F/m we can do this \[V_{f}=V _{i} + (\frac{ F }{ m }) t\] You know initial velocity is v=0 then the equation becomes \[V _{f}= (\frac{ F }{ m})t\]

OpenStudy (snowsurf):

From here its plugging in the values you are given.

OpenStudy (anonymous):

so I use the last equation correct? @snowsurf

OpenStudy (snowsurf):

Yes